### v2

@@ Acknowledgements
This work was partly supported by SoBigData.it which receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021.; and by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; and by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006); and by the ERC project DEDUCE (Data-driven and User-centered Content Moderation) under grant #101113826; and by the HORIZON Europe projects TANGO - Grant Agreement n. 101120763; and by the European Commission under the NextGeneration EU programme – National Recovery and Resilience Plan (PNRR), under agreements: PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - "FAIR - Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI".
Appendix A Heatmap overlapping matrix of the communities
(a) RTW vs. INDI.
(b) MULTI vs. INDI.
Figure A.1: Heatmap of the overlapping matrix RTW and MULTI vs. INDI.
(a) UNFL (nw) vs. INDI.
(b) UNFL (ec) vs. INDI.
(c) UNFL (sum) vs. INDI.
Figure A.2: Heatmap of the overlapping matrix UNFL vs. INDI.
[…]

### v3

@@ Acknowledgements
This work was partly supported by SoBigData.it which receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021.; and by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; and by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006); and by the ERC project DEDUCE (Data-driven and User-centered Content Moderation) under grant #101113826; and by the HORIZON Europe projects TANGO - Grant Agreement n. 101120763; and by the European Commission under the NextGeneration EU programme – National Recovery and Resilience Plan (PNRR), under agreements: PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - "FAIR - Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI"; and by eSSENCE, an e-Science collaboration funded as a strategic research area of Sweden.
CRediT authorship contribution statement
Lorenzo Mannocci: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. Stefano Cresci: Conceptualization, Methodology, Supervision, Writing – original draft, Writing – review & editing. Matteo Magnani: Conceptualization, Methodology, Supervision, Writing – review & editing. Anna Monreale: Conceptualization, Methodology, Supervision, Writing – review & editing . Maurizio Tesconi: Conceptualization, Funding acquisition, Methodology, Project administration, Resources, Supervision, Writing – review & editing.
Appendix A Additional results
In this section, we present all results for the Infomap approaches. We also report the characterization of common communities, including the starplots (Figure A.5) and the PCA visualization (Figure A.7) obtained with Louvain on IORussia.
Figure A.1: Normalized Mutual Information (NMI) heatmap comparing the communities detected with the Infomap algorithm, considering only those with size greater than 200 (for UK) and 70 (for IORussia), across all pairs of co-actions.
Figure A.2: Bar charts for Infomap approach showing the number of the lost, common, and gained communities and nodes switching from INDE to RTW in IORussia dataset.
Figure A.3: Cosine similarity between the vectors of the communities detected with Infomap in common with RTW communities. The horizontal lines report the average cosine similarity for each layer’s communities.
Figure A.4: Example of metrics comparison of three couple of common communities between RTW and MEN, HST, and URL in UK dataset with Infomap.
Figure A.5: Example of metrics comparison of three couple of common communities between RTW and MEN, HST, and URL in IORussia dataset with Louvain and Infomap.
Figure A.6: Visualization, for UK with Infomap approach of common communities between RTW and other co-actions in a two-dimensional space using PCA. Each community is represented by a vector of metrics, i.e., size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These high dimensional vectors are projected into a two-dimensional space using PCA. Communities are represented with distinct point shapes based on the co-action and are linked to their corresponding RTW community. Each RTW community may be associated with multiple communities from different co-actions, with all connected communities sharing the same color for clarity.
Figure A.7: Visualization for IORussia with Infomap and Louvain, of common communities between RTW and other co-actions in a two-dimensional space using PCA. Each community is represented by a vector of metrics, i.e., size, density, average degree, average weight, average clustering coefficient, conductance, and assortativity. These high dimensional vectors are projected into a two-dimensional space using PCA. Communities are represented with distinct point shapes based on the co-action and are linked to their corresponding RTW community. Each RTW community may be associated with multiple communities from different co-actions, with all connected communities sharing the same color for clarity.
Figure A.8: Bar charts showing the number of lost, common, and gained communities switching from INDE to UNFL (sum) and MULTI (in UK and IORussia dataset).
Figure A.9: Bar charts showing the number of lost, common, and gained nodes switching from INDE to UNFL (sum) and MULTI (in UK and IORussia dataset).
Figure A.10: Lost, common, gained nodes characterization for UNFL and MULTI (Infomap) in UK dataset. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
Figure A.11: Lost, common, gained nodes characterization for UNFL and MULTI (Louvain) in IORussia dataset. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
Figure A.12: Lost, common, gained nodes characterization for UNFL and MULTI (Infomap) in IORussia dataset. We use the notation ns and * to indicate statistical significance levels. Specifically, ns denotes that there is no significant difference between the groups, with a p-value in the range $p\geq 0.05$. For statistically significant results, we used the following annotations based on p-value ranges: *: $p\leq 0.05$; **: $p\leq 0.01$; ***: $p\leq 0.001$.
Figure A.13: Level of coordination of lost, common, and gained communities detected by Infomap approach.
Appendix B UK heatmap overlapping matrix of Louvain communities
We report the heatmaps of the overlapping matrix between Louvain communities for UK. Specifically between RTW and INDE, MULTI and INDE, UNFL (all 3 flattening strategies) and INDE.
Figure B.1: RTW vs. INDE.
Figure B.2: MULTI vs. INDE.
Figure B.3: UNFL (nw) vs. INDE.
Figure B.4: UNFL (ec) vs. INDE.
Figure B.5: UNFL (sum) vs. INDE.
Appendix C UK heatmap overlapping matrix of Infomap communities
We report the heatmaps of the overlapping matrix between Infomap communities for UK. Specifically between RTW and INDE, MULTI and INDE, UNFL (all 3 flattening strategies) and INDE.
Figure C.1: RTW vs. INDE.
Figure C.2: MULTI vs. INDE.
Figure C.3: UNFL (nw) vs. INDE.
Figure C.4: UNFL (ec) vs. INDE.
Figure C.5: UNFL (sum) vs. INDE.
Appendix D IORussia heatmap overlapping matrix of Louvain communities
We report the heatmaps of the overlapping matrix between Louvain communities for IORussia. Specifically between RTW and INDE, MULTI and INDE, UNFL (all 3 flattening strategies) and INDE.
Figure D.1: RTW vs. INDE.
Figure D.2: MULTI vs. INDE.
Figure D.3: UNFL (nw) vs. INDE.
Figure D.4: UNFL (ec) vs. INDE.
Figure D.5: UNFL (sum) vs. INDE.
Appendix E IORussia heatmap overlapping matrix of Infomap communities
We report the heatmaps of the overlapping matrix between Infomap communities for IORussia. Specifically between RTW and INDE, MULTI and INDE, UNFL (all 3 flattening strategies) and INDE.
Figure E.1: RTW vs. INDE.
Figure E.2: MULTI vs. INDE.
Figure E.3: UNFL (nw) vs. INDE.
Figure E.4: UNFL (ec) vs. INDE.
Figure E.5: UNFL (sum) vs. INDE.